TiO2 and BiVO4 are used in the ratio .2:10 the calculate the amount of each used in grams?
1
Expert's answer
2017-04-28T14:39:08-0400
If the total amount of this mixture is 100 g, than:
m (TiO2) = 100/12 ·2 = 16.7 g
m (BiVO4) = 100/12 ·10 = 83.3 g